### v1

# Real-time 3D semantic occupancy prediction for autonomous vehicles using memory-efficient sparse convolution

Abstract
In autonomous vehicles, understanding the surrounding 3D environment of the ego vehicle in real-time is essential. A compact way to represent scenes while encoding geometric distances and semantic object information is via 3D semantic occupancy maps. State of the art 3D mapping methods leverage transformers with cross-attention mechanisms to elevate 2D vision-centric camera features into the 3D domain. However, these methods encounter significant challenges in real-time applications due to their high computational demands during inference. This limitation is particularly problematic in autonomous vehicles, where GPU resources must be shared with other tasks such as localization and planning. In this paper, we introduce an approach that extracts features from front-view 2D camera images and LiDAR scans, then employs a sparse convolution network (Minkowski Engine), for 3D semantic occupancy prediction. Given that outdoor scenes in autonomous driving scenarios are inherently sparse, the utilization of sparse convolution is particularly apt. By jointly solving the problems of 3D scene completion of sparse scenes and 3D semantic segmentation, we provide a more efficient learning framework suitable for real-time applications in autonomous vehicles. We also demonstrate competitive accuracy on the nuScenes dataset.
I INTRODUCTION
The rise of Autonomous Vehicles (AV) heralds a transformative era in transportation, moving us to potentially more efficient and safer alternatives. At the heart of these autonomous systems is a software stack composed of perception, localization, planning, and control. Among these, perception is crucial as it enables the vehicle to interpret and understand its surroundings, serving as the foundation for all subsequent vehicle decision-making. Typically, AV’s deploy a variety of sensors such as cameras and LiDARs in order to perceive its surroundings. Cameras operate in 2D perspective image space, capturing essential color and texture details but lacks 3D spatial information. This contrasts with sensors like LiDAR that inherently provide 3D coordinates, albeit often with lower spatial resolution. For a holistic 3D scene understanding, both pixel-level information from cameras and sparse spatial information from LiDARs should be leveraged.
Figure 1: Scene understanding methods using camera and LiDAR in nuScenes. Top two images shows perceived sensor output. Bottom two images shows scene predictions made from sensor output.
Traditional scene understanding uses 3D object detection methods that take in sensor information from camera and LiDARs and project 3D bounding boxes across objects of interest such as vehicles, pedestrians and traffic light. However, while 3D object detection provides information about object instances, it often fails to capture the complex object geometries and background information such as drivable space and buildings [1, 2]. An alternative approach to scene understanding is constructing 3D semantic occupancy maps. Given a physical distance from the ego vehicle in length, width and height, 3D semantic occupancy map partitions the environment into a structured grid map with a predefined resolution, allowing each grid cell to be assigned a semantic label.
Nevertheless, 3D semantic occupancy prediction of outdoor scenes proves to be difficult due to the scenes’ sparsity. Large areas such as the sky and occluded space captured by the camera often contain no relevant information, and most autonomous vehicle uses rotary beam LiDARs which are inherently sparse, leading to an overall sparse data. To accurately interpret and fill in the missing parts of the environment, these algorithms must incorporate elements of 3D scene completion in addition to 3D semantic segmentation.
Thereby, our main contributions are listed as the following:
  • We propose the design of a novel sparse 3D convolution (Minkowski Engine [3]) model to perform 3D semantic occupancy prediction that jointly solves the problem of scene completion and semantic segmentation.
  • We evaluate the model’s 3D scene completion and semantic segmentation performance, achieving competitive accuracy against benchmark datasets [2, 1].
  • We conduct time and memory usage evaluations to ensure model’s real-time inference capabilities are close to human perception rates of 20 - 30 frames per second (FPS).
II Related Work
II-A Camera View Transformation
Camera view transformation techniques can be broadly bifurcated into two categories: extending 2D camera representations into a 3D domain or initializing a 3D domain and distilling 2D information onto it. The former aims to ”lift” the camera image from perspective view to a 3D view, whereas the latter aims to ”pull” camera information onto a predefined 3D space. Past lifting techniques includes probabilistic depth predictions for each camera pixel [4, 5, 6, 7, 8], as well as Multi-layer Perceptron (MLP) methods [9, 10]. A current popular pulling technique is using cross-attention based transformers where queries relayed from a discretized 3D space draws out keys and queries from a 2D perspective image[11, 12, 13, 14, 15].
Simple-BEV [16] suggests that the accuracy of Bird’s Eye View (BEV) and similarly, 3D prediction, often hinges more on factors like batch size, image resolution, and sensor fusion rather than sophisticated view transformation algorithms. Moreover, it is important to acknowledge that transforming views from 2D to 3D is inherently an ill-posed problem due to the absence of depth information in 2D images. This limitation highlights the necessity of employing deterministic methods that rely on additional sensors such as LiDARs, which are commonly included in most autonomous vehicles anyways. Arguing against the use of tools like LiDAR due to their cost overlooks their essential role in providing a fast and convenient method for accurate view transformation.
II-B 3D Semantic Scene Completion
Developments in 3D semantic scene completion of outdoor self-driving scenes is mainly spurred by well-curated datasets. With Semantic-KITTI [17], many LiDAR-based semantic scene completion methods appeared [18, 19, 20]. Given LiDAR as the only input, many of these methods do not require view transformation, simplifying the problem. However, most mainstream methods are built upon indoor LiDAR dataset such as NYUv2 [21] and ScanNet[22], and do not translate well to outdoor scenarios due to the sparsity of LiDAR points. Specific methods built upon Semantic-KITTI employed innovative solutions to densify the 3D scene. Notably, S3CNet [18] also uses Sparse convolution to balance computational budget while achieving state-of-the-art results. Nevertheless, LiDAR data lacks color and texture information, limiting its ability to discern objects with similar shapes but different visual features, such as distinguishing a street pole from a traffic light. Hence, LiDAR-only methods still require some level of feature engineering such as spherical projection to a range image, creating truncated signed distance function (TSDF) volumes, and estimating surface normal. These feature engineering inevitably requires additional tuning while introducing some level of noise.
More recently, multi-modal datasets such as nuScenes [23] and Occupancy 3D (Occ3D) [1] have enabled the development of vision-centric 3D semantic scene completion, primarily using cameras as the model input. Vision-centric methods, initially explored by Tesla [24] and further researched in various research studies [11, 1], employs cross-attention to geometrically relate 2D camera features into a 3D voxel grid, which is then used for BEV predictions or 3D semantic occupancy predictions. Efforts like deformable attention [14] and coarse-to-fine strategies [1] have been implemented to reduce the computational burden. Yet, a significant downside remains in the inherent dimensionality of establishing a dense 3D voxel grid, where attention mechanisms and transformers struggle to manage efficiently. Additionally, relying solely on camera data requires the model to simultaneously address the task of identifying and denoising distorted occluded areas on top of scene completion and semantic segmentation, posing a complex and challenging training task.
II-C Sparse Convolution
For 3D and higher dimension data, it is often inefficient to parse such data through traditional dense 3D convolutions as most grid cells are empty to begin with. In the context of 3D semantic occupancy prediction, sparse convolution operates on spatially sparse 3D data. It only considers 3D points that are specified, often ones which contains LiDAR or camera information, whilst discarding meaningless empty 3D points. Sparse Convolution Networks, like Minkowski Engine [3], have been effectively utilized in indoor scenes [25] and for single object analysis [26]. However, their application in outdoor driving scenes, particularly with the availability of multi-modal outdoor datasets, remains an area for exploration. Given their proficiency in selectively processing meaningful 3D points, these networks are well-suited for the large and sparse areas typical of LiDAR and camera data capturing outdoor environments. Adapting sparse convolution networks to outdoor 3D semantic occupancy prediction presents a promising avenue for addressing the computational challenges currently faced in this field.
III Problem Formulation
Formally, we define the 3D semantic occupancy prediction within the context of one front-view camera positioned to view directly ahead of an autonomous vehicle. Given a 2D camera image of dimensions $W\times H$, where each pixel $i_{ij}$ contains RGB values $(R,G,B)$ , the objective is to use view transformation to project and sparse convolution to densify and segment RGB and higher level image features into a 3D voxel space $V$ with dimensions $X\times Y\times Z$ in real-time. We define real-time to be at least 20 frames per second (20 FPS), with the goal of reaching 30 FPS to match human perception rate. Each voxel $v_{xyz}$ in this 3D space should encapsulate semantic and spatial information based on camera and LiDAR. The expected outcome is a labeled 3D voxel grid, where each voxel is affiliated with one of $L$ class labels, denoted as $L=\{l_{1},l_{2},...,l_{C}\}$, with $C$ representing the aggregate number of distinct classes.
IV Method
IV-A System Pipeline
The system pipeline is shown in Figure 2. Using a calibrated extrinsic matrix, $E_{\text{calib}}$ we first project LIDAR points onto the RGB image space. Each projected point is fused with RGB color information through bilinear interpolation. Subsequently, the fused data is unprojected back into the 3D space using the intrinsic camera parameters and the inverse of $E_{\text{calib}}$. We also perform feature extraction using EfficientNetV2 [27] at layer 3,4 and 6 to extract higher level 2D camera features and lift them to the 3D space using a the same projection method but scaled accordingly to accommodate for a reduction in spatial dimensions. The deterministic nature of the LiDAR-to-camera projection, assuming the sensors are adeptly calibrated, offers an immediate and precise depth-location correspondence - a key attribute for real-time tasks in autonomous vehicles.
Figure 2: Overview of System Pipeline
Initial representation now has attributes of spatial coordinates ($x,y,z$) and feature values ($[feats]+intensity$). Spatial coordinates is voxelized by a resolution of 0.4m, discretizing the pointcloud into a 3D tensor, $T_{\text{sparse}}$. $T_{\text{sparse}}$ has dimensions of 200 x 200 x 16 grid units in tensor shape, covering a physical boundary of [-40m, -40m, -1m] to [40m, 40m, 5.4m]. $T_{\text{sparse}}$ is later converted to Coordinate List format (COO) with only non-empty grids included during sparse convolution operations. We leverage the Minkowski Engine generative completion algorithm [28] to densify the sparse tensor: $T_{\text{dense}}=\mathcal{M}(T_{\text{sparse}})$, where $\mathcal{M}$ represents the generative scene completion operation. Scene completion is guided by Occ3D [1] dataset ground truth labels with the grid resolution to match $T_{\text{sparse}}$’s dimension.
We carry out semantic segmentation on $T_{\text{dense}}$, directly supervised by ground truth semantic labels. Both the semantic segmentation and scene completion step is trained as a multi-task problem, with the semantic segmentation having an additional encoder-decoder structure of smaller size, predicting outputs over 18 semantic categories. The final 3D semantic occupancy prediction has 200 x 200 x 16 voxel grid with 0.4m resolution.
IV-B Minkowski Engine
Minkowski Engine processes sparse, high-dimensional data by focusing on non-zero data points and allowing adaptability in the kernel’s shape and size. In other words, it follows a generalized convolution [3]. First, a sparse tensor is defined as a combination of a hash-table of coordinates and their corresponding features. Specifically, this is denoted as $u=[C_{n\times d},x_{n\times m}]$, where $C_{n\times d}$ represents the coordinates in COO format and $x_{n\times m}$ corresponds to the feature vectors. The convolution operation within this framework is given by:
$x_{\text{out}}^{u}=\sum_{i\in\mathcal{N}^{D}(u,C_{\text{in}})}W_{i}\cdot x_{\text{in}}^{u+i}\quad\text{for}\quad u\in C_{\text{out}}$ (1)
In equation 1, $x_{\text{out}}^{u}$ is the output feature at point $u$. $\mathcal{N}^{D}(u,C_{\text{in}})$ denotes the set of offsets defining the neighborhood of input coordinates around $u$ that plays a part in the sparse convolution layer. $W_{i}$ are the learnable weights of the convolution kernel, and $x_{\text{in}}^{u+i}$ is the input feature at the position offset by $i$ from $u$.
Figure 3: Minkowski Engine Sparse Convolution Network Architecture. Scene Completion U-Net is shown in detail. MinkUNet is the Semantic Segmentation U-Net.
IV-C Network Architecture
With reference to figure 3, the proposed 3D semantic occupancy prediction model utilizes 2 U-Net-like [29] like encoder-decoder networks sequentially. Both U-Net shares similar structure, with the scene completion U-Net having deeper layers and specialized modules such as Squeeze and Excite [30], Generative Convolution Transpose and Pruning. An encoder block contains a double convolution Layer and a squeeze and excite (SE) layer. The convolution layer features batch normalization (BN) and ReLU activation to standardize the variable density of sparse data. The SE layer dynamically recalibrates channel-wise features through global average pooling, enhancing the model’s focus on key features across the input sparse data. For certain encoder blocks, image features extracted from the feature extractor are also concatenated before further downsampling.
The decoder module for scene completion is designed for generative tasks [28]. The process begins with generative transposed convolution capable of generating new coordinates with the outer product of the weight kernel and the input coordinates. A SE layer is attached after generation. At each decoder layer, a skip connection from the corresponding encoder is also added. Following that, a 3D binary classification convolution is applied to the sparse tensor which assigns a decision value to prune. The decision threshold is trainable with ground truth supervision. Redundant features are pruned using Minkowski Engine’s built in pruning module. Overall, the decoder generates, upsamples, classifies, and then prunes voxels to iteratively create accurate 3D representation at each layer depth.
Upon reaching the final decoder layer, an additional semantic segmentation U-Net head is attached for the sole purpose of semantic segmentation of the densified sparse tensor. This U-Net features convolution layers and skip connections, with bottleneck channel at 256.
IV-D Loss Function
A multi-task loss function is defined to train the 3D Occupancy prediction model. The two terms are balanced by a Lambda ($\lambda$) constant empirically set to 0.5. Experiments conducted on different $\lambda$ constants is represented in the appendix.
$\mathcal{L}_{\text{loss}}=\mathcal{L}_{\text{completion}}+\lambda\mathcal{L}_{\text{segmentation}}$ (2)
$\mathcal{L}_{\text{completion}}$ is a voxel based binary cross entropy with logits loss ($\sum_{i,j,k}\lambda_{\text{BCE}}(p_{ijk},y_{ijk})$). It is computed over the volumetric occupancy on each decoder layer. This loss also tunes the decision threshold used for pruning generated voxels.
$\mathcal{L}_{\text{segmentation}}(z,y)=-\frac{1-\beta}{1-\beta^{n_{y}}}\log\left(\frac{\exp(z_{y})}{\sum_{j=1}^{C}\exp(z_{j})}\right)$ (3)
In normal semantic segmentation settings, multi-class cross entropy loss is sufficient with $z$ being predicted logits and $y$ being the ground truth labels. However, in nuScenes dataset, the class imbalance is extreme. For example, bicycle and motorcycle class is $10^{4}$ times lower in frequency than drivable space and free class [31]. Therefore, a class-balanced loss is applied [32], where an effective number $\beta$ is used to reweight the cross entropy loss function, giving more weight to classes with fewer effective samples. Effective number is set to $\beta=0.9$. Effective samples $\beta^{n_{y}}$ is calculated by the normalized number of voxels representing each class over all ground truth voxels over the entire training dataset. Validity of the class balanced loss is evaluated in the ablation study.
V Experiments
V-A Experimental Setup
We utilize the Occupancy 3D (Occ3D) Dataset [1], an extension of the nuScenes dataset for 3D semantic occupancy benchmarking. It comprises of 28,130 Front camera RGB training images and 6019 validation images, as well as their corresponding voxelized ground truth semantic occupancy grid of size (200x200x16, 0.4m). The input LIDAR scans are obtained from the nuScenes [23] dataset. The training was configured with the Adam optimizer at a learning rate of 1e-4, and a Cosine-Annealing learning rate schedule. The model was trained till convergence, with a batch size of 10. The hardware setup for the training included 1 NVIDIA RTX 4090 GPU and 24GB DDR4 RAM. Training VRAM usage is at 4.3 GB.
Data augmentation is performed on the input and ground truth dataset, which includes RGB and LiDAR intensity value perturbation to simulate sensor noise (noise strength of 0.05 on normalized RGB values), random translation (-4 to 4 voxel grids), random yaw rotation (+/- 45 degrees), and ground truth voxel masking (5 percent).
V-B Scene Completion
We evaluate the 3D scene completion performance using binary voxel completion intersect over union (IoU), precision, recall, and F1 score to be compared with other scene completion methods. Referring to Table I, $P$ represents the set of voxels predicted by the model, $GT$ the set of voxels in the dense occupancy ground truth.
TABLE I: Evaluation metrics for 3D scene completion task.
Metric Equation Better
IoU $\frac{|P\cap GT|}{|P\cup GT|}$ Higher
Precision $\frac{|P\cap GT|}{|P|}$ Higher
Recall $\frac{|P\cap GT|}{|GT|}$ Higher
F1 Score $2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ Higher
TABLE II: Quantitative results of 16 semantic classes on nuScenes validation dataset.
mIoU
Barrier
Bicycle
Bus
Car
Cons. Veh.
Motorcycle
Pedestrian
Traf. Cone
Trailer
Truck
Driv. Sur.
Other Flat
Sidewalk
Terrain
Manmade
Vegetation
Ours (%) 36.03 40.12 0.00 51.01 63.10 12.52 3.60 30.16 7.69 32.71 32.17 87.19 37.42 41.34 44.46 74.24 84.68
BEVFormer [11] (%) 23.70 36.77 11.70 29.87 38.92 10.29 22.05 16.21 14.69 26.44 23.13 48.19 33.10 29.80 17.64 19.01 13.75
Occ3D [1] (%) 28.53 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 33.84 37.98 33.23 20.79 18.0
NVOCC [33] (%) 54.19 57.98 46.40 52.36 63.07 35.68 48.81 42.98 41.75 60.82 49.56 87.29 58.29 65.93 63.30 64.28 53.76
Figure 4: Qualitative predictions on nuScenes validation dataset. Semantic occupancy grid is viewed in 3rd person perspective angle. Each group of images consists of occupancy prediction image on the left, ground truth label from Occ3D [1] on the right, and front camera image on top.
We compare these metrics with other scene completion methods on the nuScenes validation dataset. In particular, TransformerFusion [34] and Atlas [35] originally targets the task of 3D reconstruction of indoor scenes from RGB-D images. SurroundOcc [36], MonoScene [37] and Occ3D [1] focuses on the same outdoor driving scene completion task. Values from other methods are achieved through dense ground truth supervision for fair comparison, as reported by [36]. Referring to Table III, our proposed method achieves an scene completion IoU of 0.533, outperforming the current state-of-art by 16%. We also show a large improvement in precision, recall and f1, indicating we densify much more of the sparse input scene while being accurate to the ground truth.
TABLE III: Scene Completion evaluation on nuScenes validation set.
Model IoU Precision Recall F1 Score
TransformerFusion [34] - 0.375 0.591 0.453
Atlas [35] - 0.407 0.546 0.458
SurroundOcc [36] 0.347 0.414 0.602 0.483
MonoScene [37] 0.342 - - -
Occ3D [1] 0.377 - - -
Ours 0.533 0.775 0.632 0.690
V-C Semantic Segmentation
In semantic segmentation, mIoU (mean Intersection over Union) is calculated using the following formula: $\text{mIoU}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$, where $N$ represents the total number of semantic classes, $TP_{i}$ denotes True Positives for class $i$, $FP_{i}$ represents False Positives for class $i$, and $FN_{i}$ corresponds to False Negatives for class $i$. We monitored the semantic segmentation performance through mean Intersection over Union (mIoU) and semantic IoU for each semantic class, and compare it with baseline BEVFormer [11], Occ3D [1], and NVOCC [33], current state-of-the-art. We do not compare the void and free class. Referring to Table II, we observe a 8 - 10 % improvement in mIoU over BEVFormer and Occ3D, as well as improvements in most semantic class IoU. Compared to NVOCC, we are competitive only in car, drivable surface, manmade and vegetation. However, bicycle, motorcycle and traffic cones presents a significant challenge to the model’s prediction. These small objects not only appear a lot less in Occ3D’s and nuScenes’s dataset, but also might be smoothed in the convolution processes, erasing or blending their features with its surroundings.
Figure 4 shows qualitative inferences of the model. Left group of day-time images demonstrates the model’s ability to classify scene objects of varying distance and shapes into distinct semantic classes. Right group of images shows model’s fidelity in low light environments: it is able to discern dark colored vehicles at night, as well as predicting obstacles on the left side of the drivable surface. However, the model does show a decrease in maximum prediction distance. For more qualitative results on different scenes, please refer to the appendix.
V-D Real-time Inference
Inference are performed on RTX 4090 with six multi-camera images for the other models at 1600x900 image resolution. Since our model focuses on monocular camera, we use a batch size of 6 for fair comparison. We compare the inference time and GPU memory usage in Table IV, showing a 6 - 10 times increase in inference time, translating to a potential frame rate of 30 FPS and an achievable frame rate of 20 FPS. while utilizing only 1.2GB of GPU memory. For real-time inference demo video, please refer to our supplementary material.
TABLE IV: Model time efficiency and GPU memory usage comparison
Model Inference Time (s) Memory Usage (MB)
BEVformer [11] 0.31 4500
MonoScene [37] 0.87 20300
SurroundOcc [36] 0.34 5900
Occ3D [1] 0.38 18000
Ours 0.03 - 0.05 1200
V-E Ablation Study
Referring to Table V, we conducted ablation studies by removing the class balanced cross entropy loss and use standard focal loss, removing the squeeze and excite layer, and removing the higher level feature input tensor. ”Ours” refers to the current model with all three components included. The results are validated against nuScenes validation dataset. The semantic segmentation mIoU and completion IoU are reported. We observed that removing the class balance loss results in a severe drop in both completion IoU and mIoU. This validates the problem of class imbalance in nuScenes dataset, and class balance loss correctly addresses it by leading the model to be sensitive to all classes in both scene completion and semantic segmentation. The incorporation of Squeeze-and-Excite layers leads to slight improvements in both mIoU and completion IoU. This can be attributed to the use of global average pooling, which enhances the model’s adaptability to varying outdoor scenes. Lastly, concatenating higher level image features with the initial $RGB+Intensity$ features achieved no significant performance improvements. This outcome could suggest that the initial feature set was likely sufficient for the task. It could also indicate the feature extractor, EfficientNetV2 [27], might have limited capacity to produce meaningful outdoor scene features.
TABLE V: Ablation Study on three core model components.
Description
Semseg mIoU Completion IoU
Ours
0.3603 0.5329
Ours w/o CBLoss
0.3240 0.3911
Ours w/o SELayer
0.3360 0.4983
Ours w/o features
0.3622 0.5110
VI Conclusions and future work
In this paper, we presented a Minkowski engine-based sparse convolution model capable of semantic scene completion on diverse outdoors autonomous vehicle driving scenes using LiDAR and a front-view camera. In particular, we demonstrated that the model is able to perform 3D semantic occupancy prediction with much less computational resources as compared to many current methods, achieving real-time inference, while being on par in accuracy. More broadly, the sparse convolution approach we propose for AVs offers a pathway to enhance the efficiency of real-time robotics perception.
Future work involves several aspects. Firstly, we hope to extend the current model to a multi-view camera setup for 360 degree occupancy prediction of the ego vehicle. Secondly, it is challenging to perform scene completion at distances beyond 20 meters due to the sparsity of LiDAR points (specifically, a 32-beam setup in nuScenes [23]). Incorporating probabilistic techniques to integrate distant camera features into 3D space may help achieve better accuracy. Thirdly, performing self-supervision by creating pseudo dense occupancy ground truth via monocular 2D camera-based depth estimation and semantic segmentation helps to expand the training dataset, allowing the current model to be more robust to complex driving environments. Lastly, on the issue of handling heavily occluded scenes, the integration of a generative adversarial sparse network as a discriminator into the existing scene completion pipeline can be investigated.
APPENDIX
VI-A More Qualitative Examples
Figure 5: This example shows the model’s ability to predict dynamic obstacles such as temporary road barriers installed for construction on drivable surface.
Figure 6: This example shows the model’s inability to hallucinate occluded areas caused by blocking cars. This is the expected performance as the model should not imagine areas without a proper occlusion handling method.
Figure 7: This example shows a night time intersection with multiple vehicles. The model is able to detect both vehicles with their geometries. However, night time prediction of drivable surface is poor and further distances are empty.
VI-B Lambda constant investigation
Using a lambda value of 0.5 yields the best results, with a minor yet notable improvement in Semantic Segmentation mIoU to 0.3555 and a more significant enhancement in Completion IoU to 0.5243.
TABLE VI: Lambda constant effect on 3D semantic occupancy prediction.
$\lambda$ Semseg mIoU Completion IoU
0.3 0.3534 0.5089
0.5 0.3555 0.5243
0.7 0.3466 0.5115
ACKNOWLEDGMENT
This project was supported by the EPSRC project RAILS (EP/W011344/1) and the Oxford Robotics Institute’s research project RobotCycle.